# Improved bounds for randomized Schatten norm estimation of numerically low-rank matrices

### v2

@@ 3.2.1 Breaking Var($\hat{\theta}_{2p}$) into mathematically manageable pieces
Let $A$ be an $m\times n$ deterministic matrix and $\Omega\in\mathbb{R}^{n\times k}$ a random matrix with i.i.d. $N(0,1)$ entries. Considering the full singular value decomposition $A=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$, $\Sigma\in\mathbb{R}^{m\times n}$, and $V\in\mathbb{R}^{n\times n}$, the entries of $\widetilde{\Omega}:=V^{T}\Omega$ are also i.i.d. $N(0,1)$ random variables, due to the rotation invariance of the standard Gaussian distribution. Therefore, the matrix $Z$ in line 1 of Algorithm 1 can be written as
$Z=(A\Omega)^{T}(A\Omega)=\widetilde{\Omega}^{T}\Sigma^{T}U^{T}U\Sigma\widetilde{\Omega}=\widetilde{\Omega}^{T}\Sigma^{T}\Sigma\widetilde{\Omega}.$
@@
with the convention that $\sigma_{\min\{n,m\}+1}=\ldots=\sigma_{\max\{n,m\}}=0$ and using the relation (4), the estimator $\hat{\theta}_{2p}$ from Algorithm 1 takes the form:
$\hat{\theta}_{2p}$ $=\binom{k}{p}^{-1}\sum_{i_{1}<\cdots<i_{p}}(\sigma_{1}^{2}\omega_{1,i_{1}}\omega_{1,i_{2}}+\cdots+\sigma_{n}^{2}\omega_{n,i_{1}}\omega_{n,i_{2}})\cdots(\sigma_{1}^{2}\omega_{1,i_{p}}\omega_{1,i_{1}}+\cdots+\sigma_{n}^{2}\omega_{n,i_{p}}\omega_{n,i_{1}})$
@@
with the indices $i_{1},\ldots,i_{p}$ ranging from $1$ to $k$ and with the convention that $i_{p+1}\equiv i_{1}$. For ease of notation, define $\hat{\Theta}_{2p}:=\binom{k}{p}\hat{\theta}_{2p}$ so that
@@
$\hat{\Theta}_{2p}^{2}=\sum_{i_{1}<\cdots<i_{p}\atop j_{1}<\cdots<j_{p}}\prod_{h=1}^{p}\left[\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right)\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,j_{h}}\omega_{t,j_{h+1}}\right)\right],$ (6)
where the indices $i_{1},\ldots,i_{p}$ and $j_{1},\ldots,j_{p}$ range from $1$ to $k$. Estimating or bounding the variance of $\hat{\theta}_{2p}$ involves estimating or bounding the expectation of each term in the outer sum from (6). For ease of notation, we denote
$f_{i_{1},\ldots,j_{p}}:=\prod_{h=1}^{p}\left[\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right)\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,j_{h}}\omega_{t,j_{h+1}}\right)\right]$ (7)
and let $F_{r}$ be the sum of $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ over the choices of $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r$ repeated indices:
$F_{r}:=\sum_{2p\text{-tuples }i_{1}<\cdots<i_{p},j_{1}<\cdots<j_{p}\text{ among }\atop\text{ which we have }r\text{ repeated indices}}\mathbb{E}[f_{i_{1},\ldots,j_{p}}].$ (8)
Then, we can split the second moment of $\hat{\Theta}_{2p}$ as
$\mathbb{E}[\hat{\Theta}_{2p}^{2}]=\sum_{r=0}^{p}\left(\sum_{2p\text{-tuples }i_{1}<\cdots<i_{p},j_{1}<\cdots<j_{p}\text{ among }\atop\text{ which we have }r\text{ repeated indices}}\mathbb{E}[f_{i_{1},\ldots,j_{p}}]\right)=\sum_{r=0}^{p}F_{r}.$ (9)
Our analysis follows by estimating or bounding each $F_{r}$. Lemmas 6 and 7 below describes the contribution of each $F_{r}$ to $\mathbb{E}[\hat{\Theta}_{2p}^{2}]$ in terms of the order of $k$.
@@
The number of $2p$-tuples $i_{1},\ldots,i_{p},j_{1},\ldots,j_{p}$ such that $i_{1}<\cdots<i_{p}$, $j_{1}<\cdots<j_{p}$, and among $i_{1},\ldots,j_{p}$ there are $r$ repeated indices $(0\leq r\leq p)$ is $\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}$.
@@ 3.2.2 First-order analysis
Since $\binom{k}{p}^{2}$ is a polynomial (in $k$) of degree $2p$ and, by Lemma 7, $F_{r}$ is a polynomial in $k$ of degree $2p-r$, it suffices to consider $F_{0}$ and $F_{1}$ in the first-order analysis. Therefore, we are interested in the sum of $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ over the choices of $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with either zero or one repeated index. Expanding the product of the two summations in (7), the expectation $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ can be further decomposed into the sum of $n^{2p}$ expectations:
@@
Observe that, if any of the $\omega$-variables appearing in (12) has an odd exponent, the expectation of $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ becomes zero since all odd moments of standard Gaussian random variables are zero. This observation gives the following proposition.
Proposition 8.
@@
Leveraging Proposition 8, we can compute $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ in the cases where there are either no repeated indices or only one repeated index among $i_{1},\ldots,j_{p}$.
@@
If there is no repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
@@
Consider a single nonzero term $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ on the right-hand-side of (11). Then $t_{1}=\ldots=t_{p}$ and $s_{1}=\ldots=s_{p}$ by Proposition 8 since there is no repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$. In this case, we have
@@
If there is exactly one repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
@@
Let $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ be a nonzero term on the right-hand-side of (11). Without loss of generality, let us assume that the repeated index is $i_{1}$. Then Proposition 8 guarantees that $t_{h}=t_{h-1}$ for $h=2,\ldots,p$, so that $t_{1}=\ldots=t_{p}$. The same argument also proves $s_{1}=\ldots=s_{p}$.
@@
Given Lemmas 9 and 10, we are ready to compute the first-order expansion of $\mathrm{Var}(\hat{\theta}_{2p})$ in $1/k$ presented in Theorem 3.
@@
Combining Lemmas 6, 9 and 10, we obtain
@@ 3.2.3 Terms with at least 2 repeated indices
To obtain an upper bound for $\mathrm{Var}(\hat{\theta}_{2p})$, we need to further consider $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r$ repeated indices and $r\geq 2$. Motivated by the decomposition in (11), we essentially want to count or bound the number of tuples $(t_{1},\ldots,s_{p})\in[n]^{2p}$ such that $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ has nonzero expectation. To better visualize the product
@@
Let $i_{h}$ be a non-repeated index (corresponding to non-circled point in Figure 2). Proposition 8 asserts that the two segments $(i_{h-1},i_{h})$ and $(i_{h},i_{h+1})$ adjacent to point $i_{h}$ must be assigned to the same singular value; otherwise the expectation of $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ will vanish. To illustrate this fact, in Figure 2, we used the same color to highlight the segments that need to be assigned to the same singular value as a necessary condition for the expectation of the term (12) to be nonzero. In the following, we refer to the group of contiguous segments with the same color as an “arc”.
Consider a choice of indices $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r$ repeated indices. Then, each circle has $r$ colored arcs. Moreover, there is a natural bijection from the colored arcs in $i$-circle to the color arcs in $j$-circle, where each colored arcs is mapped to the colored arc in the other circle with the same indices as endpoints. In our example, the pairings are $(i_{1},i_{2})$$(j_{2},j_{4})$, $(i_{2},i_{5})$$(j_{4},j_{8})$, and $(i_{5},i_{1})$$(j_{8},j_{2})$, where each arc is intended in a counter-clockwise direction. We denote the arcs by $a_{1}^{(i)},\ldots,a_{r}^{(i)},a_{1}^{(j)},\ldots,a_{r}^{(j)}$ and we will say that $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ are corresponding pairs of arcs for $\ell=1,\ldots,r$.
The next proposition limits the possible singular values assignment to $2r$ arcs in the two circles and is crucial for the proofs of Lemma 13 and Lemma 14.
Proposition 11.
Fix a choice of $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r\geq 2$ repeated indices and a $2p$-tuple $(t_{1},\ldots,s_{p})\in[n]^{2p}$ such that $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ is nonzero. Consider the visualization for $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ as in Figure 2. Suppose the arcs $a_{1}^{(i)}$ and $a_{2}^{(i)}$ in the $i$-circle are assigned to singular values $\sigma_{g}$ and $\sigma_{h}$, respectively, with $g\neq h$. Then there are only two possible singular values assignments to their corresponding arcs $a_{1}^{(j)}$ and $a_{2}^{(j)}$ in the $j$-circle:
@@
Now, we are ready to bound $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ when there are two or more repeated indices among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$.
@@
If there are exactly two repeated indices among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
@@
Since there are two repeated indices, there are two arcs for each circle. Using our previous notation, they are $a_{1}^{(i)}$, $a_{2}^{(i)}$, $a_{1}^{(j)}$, and $a_{2}^{(j)}$. By Proposition 11, the nonzero expectation $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ on the right-hand-side of (11) corresponds to one of the following situations.
@@
If there are $r>2$ repeated indices among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$, then
@@
It suffices to consider the terms $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ with nonzero expectation. By Proposition 11, the maximum number of different colors in the two circles of Figure 2 (i.e., the maximum number of different singular values in the product (14)) is $r$. We separately consider the terms $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ in which there are $\ell$ distinct colors (singular values), for $1\leq\ell\leq r$.
@@
  • Case $\ell=2$. Without loss of generality, assume the two colors are $t_{1}$ and $t_{2}$. First of all, the maximum number of fourth powers of random variables in the product (14) is $r-2$, which gives a coefficient of at most $3^{r-2}$ in front of the singular values. Due to Proposition 11, there are only two possible cases 1 and 2:
@@
  • Case $3\leq\ell\leq r$. Denote the distinct colors by $t_{1},\ldots,t_{\ell}$. First of all, by Proposition 11, only case 1 can happen: this means that, for each $\ell=1,\ldots,r$, the two corresponding arcs $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ have the same color.
@@
    Since we are under Proposition 11, case 1, the color of $a_{\ell}^{(i)}$ uniquely determines the color of $a_{\ell}^{(j)}$. It therefore suffices to determine the color pattern for $i$-circle. A valid color pattern needs to have all $\ell$ different colors in $i$-cycle. The total number of valid color patterns is bounded by $\ell^{r}$ (which is the number of ways to assign $\ell$ colors to $r$ arcs in $i$-circle).

### v3

@@ 3.2.1 Breaking Var($\hat{\theta}_{2p}$) into mathematically manageable pieces
Let $A$ be an $m\times n$ deterministic matrix and $\Omega\in\mathbb{R}^{n\times k}$ a random matrix with i.i.d. $N(0,1)$ entries. Considering the full singular value decomposition $A=U\Sigma V^{T}$, where $U\in\mathbb{R}^{m\times m}$, $\Sigma\in\mathbb{R}^{m\times n}$, and $V\in\mathbb{R}^{n\times n}$, the entries of $\widetilde{\Omega}:=V^{T}\Omega$ are also i.i.d. $N(0,1)$ random variables, due to the rotation invariance of the standard Gaussian distribution. Then
$\Omega^{T}A^{T}A\Omega=\widetilde{\Omega}^{T}\Sigma^{T}U^{T}U\Sigma\widetilde{\Omega}=\widetilde{\Omega}^{T}\Sigma^{T}\Sigma\widetilde{\Omega}.$
@@
with the convention that $\sigma_{\min\{n,m\}+1}=\ldots=\sigma_{\max\{n,m\}}=0$ and using the first identity in (4), the estimator $\hat{\theta}_{2p}$ from Algorithm 1 takes the form:
$\hat{\theta}_{2p}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\binom{k}{p}^{-1}\sum_{i_{1}<\cdots<i_{p}}(\widetilde{\Omega}^{T}\Sigma^{T}\Sigma\widetilde{\Omega})_{i_{1}i_{2}}(\widetilde{\Omega}^{T}\Sigma^{T}\Sigma\widetilde{\Omega})_{i_{2}i_{3}}\cdots(\widetilde{\Omega}^{T}\Sigma^{T}\Sigma\widetilde{\Omega})_{i_{p}i_{1}}}$
$=\binom{k}{p}^{-1}\sum_{i_{1}<\cdots<i_{p}}(\sigma_{1}^{2}\omega_{1,i_{1}}\omega_{1,i_{2}}+\cdots+\sigma_{n}^{2}\omega_{n,i_{1}}\omega_{n,i_{2}})\cdots(\sigma_{1}^{2}\omega_{1,i_{p}}\omega_{1,i_{1}}+\cdots+\sigma_{n}^{2}\omega_{n,i_{p}}\omega_{n,i_{1}})$
@@
with the indices $i_{1},\ldots,i_{p}$ ranging from $1$ to $k$ and with the convention that $i_{p+1}\equiv i_{1}$. For ease of notation, we work with the rescaled quantity $\hat{\Theta}_{2p}$ defined in (5). Then,
@@
$\hat{\Theta}_{2p}^{2}=\sum_{i_{1}<\cdots<i_{p}\atop j_{1}<\cdots<j_{p}}\prod_{h=1}^{p}\left[\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right)\left(\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}=1}^{n}\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{2}\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s},j_{h}}\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s},j_{h+1}}\right)\right],$ (8)
where the indices $i_{1},\ldots,i_{p}$ and $j_{1},\ldots,j_{p}$ range from $1$ to $k$. Estimating or bounding the variance of ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\hat{\Theta}_{2p}}$ involves estimating or bounding the expectation of each term in the outer sum from (8). For ease of notation, we denote
$f_{i_{1},\ldots,j_{p}}:=\prod_{h=1}^{p}\left[\left(\sum_{t=1}^{n}\sigma_{t}^{2}\omega_{t,i_{h}}\omega_{t,i_{h+1}}\right)\left(\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}=1}^{n}\sigma_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s}}^{2}\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s},j_{h}}\omega_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}s},j_{h+1}}\right)\right];$ (9)
we define
$\mathcal{I}_{r}:=\left\{2p\text{-tuples }i_{1}<\cdots<i_{p},j_{1}<\cdots<j_{p}\text{ among which we have }r\text{ repeated indices}\right\}$
and let
$F_{r}:=\sum_{(i_{1},\ldots,j_{p})\in\mathcal{I}_{r}}\mathbb{E}[f_{i_{1},\ldots,j_{p}}].$ (10)
be the sum of $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ over the choices of $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ with $r$ repeated indices. Then, we can split the second moment of $\hat{\Theta}_{2p}$ as
$\mathbb{E}[\hat{\Theta}_{2p}^{2}]=\sum_{r=0}^{p}\left(\sum_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(i_{1},\ldots,j_{p})\in\mathcal{I}_{r}}}\mathbb{E}[f_{i_{1},\ldots,j_{p}}]\right)=\sum_{r=0}^{p}F_{r}.$ (11)
Our analysis follows by estimating or bounding each $F_{r}$. Lemmas 9 and 10 below describes the contribution of each $F_{r}$ to $\mathbb{E}[\hat{\Theta}_{2p}^{2}]$ in terms of the order of $k$.
@@
The cardinality of the set $\mathcal{I}_{r}$, for $0\leq r\leq p$, is $\binom{k}{2p-r}\binom{2p-r}{p}\binom{p}{r}$.
@@ 3.2.2 First-order analysis
Since $\binom{k}{p}^{2}$ is a polynomial (in $k$) of degree $2p$ and, by Lemma 10, $F_{r}$ is a polynomial in $k$ of degree $2p-r$, it suffices to consider $F_{0}$ and $F_{1}$ in the first-order analysis. Therefore, we are interested in the sum of $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ over the choices of $2p$-tuples $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$ in either $\mathcal{I}_{0}$ or $\mathcal{I}_{1}$. Expanding the product of the two summations in (9), the expectation $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ can be further decomposed into the sum of $n^{2p}$ expectations:
@@
Observe that, if any of the $\omega$-variables appearing in (14) has an odd exponent, the expectation of $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ becomes zero since all odd moments of standard Gaussian random variables are zero. This observation gives the following lemma.
Lemma 11.
@@
Leveraging Lemma 11, we can compute $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ in the cases where there are either no repeated indices or only one repeated index among $i_{1},\ldots,j_{p}$.
@@
If $(i_{1},\ldots,j_{p})\in\mathcal{I}_{0}$ then
@@
Consider a single nonzero term $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ on the right-hand-side of (13). Then $t_{1}=\ldots=t_{p}$ and $s_{1}=\ldots=s_{p}$ by Lemma 11 since there is no repeated index among $i_{1}<\cdots<i_{p}$ and $j_{1}<\cdots<j_{p}$. In this case, we have
@@
If $(i_{1},\ldots,j_{p})\in\mathcal{I}_{1}$ then
@@
Let $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ be a nonzero term on the right-hand-side of (13). Without loss of generality, let us assume that the repeated index is $i_{1}$. Then Lemma 11 guarantees that $t_{h}=t_{h-1}$ for $h=2,\ldots,p$, so that $t_{1}=\ldots=t_{p}$. The same argument also proves $s_{1}=\ldots=s_{p}$.
@@
Given Lemmas 12 and 13, we are ready to compute the first-order expansion of $\mathrm{Var}(\hat{\theta}_{2p})$ in $1/k$ presented in Theorem 2.
@@
Combining Lemmas 9, 12 and 13, we obtain
@@ 3.2.3 Terms with at least 2 repeated indices
To obtain an upper bound for $\mathrm{Var}(\hat{\theta}_{2p})$, we need to further consider $2p$-tuples $(i_{1},\ldots,j_{p})\in\mathcal{I}_{r}$ with $r\geq 2$. Motivated by the decomposition in (13), we essentially want to count or bound the number of tuples $(t_{1},\ldots,s_{p})\in[n]^{2p}$ such that $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ has nonzero expectation. To better visualize the product
@@
Let $i_{h}$ be a non-repeated index (corresponding to non-circled point in Figure 2). Lemma 11 asserts that the two segments $(i_{h-1},i_{h})$ and $(i_{h},i_{h+1})$ adjacent to point $i_{h}$ must be assigned to the same singular value; otherwise the expectation of $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ will vanish. To illustrate this fact, in Figure 2, we used the same color to highlight the segments that need to be assigned to the same singular value as a necessary condition for the expectation of the term (14) to be nonzero. In the following, we refer to the group of contiguous segments with the same color as an “arc”.
Consider a $2p$-tuple $(i_{1},\ldots,j_{p})\in\mathcal{I}_{r}$. Then, each circle has $r$ colored arcs. Moreover, there is a natural bijection from the colored arcs in $i$-circle to the color arcs in $j$-circle, where each colored arcs is mapped to the colored arc in the other circle with the same indices as endpoints. In our example, the pairings are $(i_{1},i_{2})$$(j_{2},j_{4})$, $(i_{2},i_{5})$$(j_{4},j_{8})$, and $(i_{5},i_{1})$$(j_{8},j_{2})$, where each arc is intended in a counter-clockwise direction. We denote the arcs by $a_{1}^{(i)},\ldots,a_{r}^{(i)},a_{1}^{(j)},\ldots,a_{r}^{(j)}$ and we will say that $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ are corresponding pairs of arcs for $\ell=1,\ldots,r$.
The next lemma limits the possible singular values assignment to $2r$ arcs in the two circles and is crucial for the proofs of Lemma 16 and Lemma 17.
Lemma 14.
Fix a $2p$-tuple $(i_{1},\ldots,j_{p})\in\mathcal{I}_{r}$ and a $2p$-tuple $(t_{1},\ldots,s_{p})\in[n]^{2p}$ such that $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ is nonzero. Consider the visualization for $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ as in Figure 2. Suppose the arcs $a_{1}^{(i)}$ and $a_{2}^{(i)}$ in the $i$-circle are assigned to singular values $\sigma_{g}$ and $\sigma_{h}$, respectively, with $g\neq h$. Then there are only two possible singular values assignments to their corresponding arcs $a_{1}^{(j)}$ and $a_{2}^{(j)}$ in the $j$-circle:
@@
Now, we are ready to bound $\mathbb{E}[f_{i_{1},\ldots,j_{p}}]$ when the $2p$-tuple $(i_{1},\ldots,j_{p})\in\mathcal{I}_{r}$ with $r\geq 2$.
@@
If $(i_{1},\ldots,j_{p})\in\mathcal{I}_{2}$, then
@@
Since there are two repeated indices, there are two arcs for each circle. Using our previous notation, they are $a_{1}^{(i)}$, $a_{2}^{(i)}$, $a_{1}^{(j)}$, and $a_{2}^{(j)}$. By Lemma 14, the nonzero expectation $\mathbb{E}[Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}]$ on the right-hand-side of (13) corresponds to one of the following situations.
@@
If $(i_{1},\ldots,j_{p})\in\mathcal{I}_{r}$ for some $r>2$, then
@@
It suffices to consider the terms $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ with nonzero expectation. By Lemma 14, the maximum number of different colors in the two circles of Figure 2 (i.e., the maximum number of different singular values in the product (16)) is $r$. We separately consider the terms $Y_{i_{1},\ldots,j_{p}}^{t_{1},\ldots,s_{p}}$ in which there are $\ell$ distinct colors (singular values), for $1\leq\ell\leq r$.
@@
  • Case $\ell=2$. Without loss of generality, assume the two colors are $t_{1}$ and $t_{2}$. First of all, the maximum number of fourth powers of random variables in the product (16) is $r-2$, which gives a coefficient of at most $3^{r-2}$ in front of the singular values. Due to Lemma 14, there are only two possible cases 1 and 2:
@@
  • Case $3\leq\ell\leq r$. Denote the distinct colors by $t_{1},\ldots,t_{\ell}$. First of all, by Lemma 14, only case 1 can happen: this means that, for each $\ell=1,\ldots,r$, the two corresponding arcs $a_{\ell}^{(i)}$ and $a_{\ell}^{(j)}$ have the same color.
@@
    Since we are under Lemma 14, case 1, the color of $a_{\ell}^{(i)}$ uniquely determines the color of $a_{\ell}^{(j)}$. It therefore suffices to determine the color pattern for $i$-circle. A valid color pattern needs to have all $\ell$ different colors in $i$-cycle. The total number of valid color patterns is bounded by $\ell^{r}$ (which is the number of ways to assign $\ell$ colors to $r$ arcs in $i$-circle).